OpenStudy (anonymous):

find the exact value of sin 17pi/12 using sum and difference formulas

terenzreignz (terenzreignz):

Okay, first, let's review... What is \[\Large \sin(\pi+\alpha)\]?

OpenStudy (anonymous):

well from what i learned sin(x+y)=sinxcosy+cosxsiny, im not quite sure what your talking about.

terenzreignz (terenzreignz):

Yeah, using that formula, we do, indeed, get... \[\Large \sin\pi\cos\alpha+\cos\pi\sin\alpha\]

terenzreignz (terenzreignz):

So, what's sin π? What's cos π?

OpenStudy (anonymous):

okay, so what i did was split 17pi/12 into sin(8pi/12+9pi/12)=sin(4pi/3+3pi/4) but im not sure what to do from there.

OpenStudy (anonymous):

ummm i dont know

terenzreignz (terenzreignz):

Well, that works. Now, there's a good time to apply your formula. By the way, sin π = 0 cos π = -1 Anyway, apply your formula for \[\Large \sin\left(\frac{8\pi}{12}+\frac{9\pi}{12}\right)\] Also, you might want to re-examine your reduction to lowest terms.

terenzreignz (terenzreignz):

I think you'll find that \[\Large \frac{8\pi}{12}=\frac{2\pi}{3}\ne\frac{4\pi}{3}\]

OpenStudy (anonymous):

oh okay so if i apply every thing i got sin(2pi/3)*cos(3pi/4)+cos(2pi/3)*sin(3pi/4) right?

terenzreignz (terenzreignz):

That's right :) Have you got those values down?

OpenStudy (anonymous):

like from a unit circle u mean?

terenzreignz (terenzreignz):

From wherever, as long as you got them :)

terenzreignz (terenzreignz):

But yeah, the unit circle ^.^

OpenStudy (anonymous):

okay um 2pi/3= 120 degrees or (-1/2, sqrt3/2) and 3pi/4= 135 degrees or -(sqrt2/2, sqrt2/2)

terenzreignz (terenzreignz):

Yeah... so, you got your value down?

OpenStudy (anonymous):

okay on a calculator the answer is -(sqrt6/4) -(sqrt2/4)

OpenStudy (anonymous):

is that right?

terenzreignz (terenzreignz):

The calculator's not wrong... but did you arrive at that answer yourself?

OpenStudy (anonymous):

nope, i dont know how to solve it out..

OpenStudy (anonymous):

but thanks for your help anyways :)

terenzreignz (terenzreignz):

@sevenupsevendown Do you want to arrive at it manually (ie, no calculators) ?

OpenStudy (anonymous):

yes please

terenzreignz (terenzreignz):

Well, by your formula, \[\large \sin\left(\frac{2\pi}{3}+\frac{3\pi}{4}\right)=\sin\frac{2\pi}{3}\cos\frac{3\pi}{4}+\cos\frac{2\pi}{3}\sin\frac{3\pi}{4}\]

OpenStudy (anonymous):

ahuh i got that part then what?

terenzreignz (terenzreignz):

Now, on the unit circle, \[\huge C(\theta)=(\cos\theta,\sin\theta)\] You got this part as well?

terenzreignz (terenzreignz):

So, \[\huge C\left(\frac{2\pi}{3}\right)=C(120^o)=\left(-\frac12 \ , \ \frac{\sqrt3}2\right)\]

terenzreignz (terenzreignz):

So, \[\Large \cos\left(\frac{2\pi}{3}\right)=-\frac12\]\[\Large\sin\left(\frac{2\pi}{3}\right)=\frac{\sqrt3}2\]

terenzreignz (terenzreignz):

Make your presence felt. ^.^

OpenStudy (anonymous):

how did u get that last part?

terenzreignz (terenzreignz):

\[\huge C(\theta)=(\cos\theta,\sin\theta)\]

terenzreignz (terenzreignz):

 Unit circle

terenzreignz (terenzreignz):



OpenStudy (anonymous):

oooooooooh i get it :)

terenzreignz (terenzreignz):

Now, can you tell me what \[\Large \cos\left(\frac{3\pi}{4}\right)\]and\[\Large \sin\left(\frac{3\pi}{4}\right)\]are?

OpenStudy (anonymous):

umm the first one is -(sqrt2/2) and the second one is (sqrt2/2)

terenzreignz (terenzreignz):

Then just compute :)

OpenStudy (anonymous):

whooo hooo! thank you!